Session: Mar 2010 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Roda at Sun Mar 21 11:09:04 2010) Failure:RECE RECO FS
A lot of prblems Problem with recorder MK5 and computer problems, lost scan from 1 to 29. For Receiver wrong setup probably error in scan from 30 to 118

Onsala (Michael Lindqvist at Sun Mar 21 11:55:01 2010) Success:
Some rain. Otherwise no known problems.

Yebes (Pablo de Vicente at Sun Mar 21 12:47:24 2010) Failure: ANT
The antenna stopped working at 22:16 UTC on previous day due to a hardware/software problem still to be diagnosed and fixed. We did not take part in the observation.

Shanghai (Bo Xia at Mon Mar 22 02:43:48 2010) Success:

Torun (K. Borkowski at Mon Mar 22 08:19:13 2010) Success:
No known problems.

Westerbork (Geert Kuper at Mon Mar 22 15:00:57 2010) Success:
No known problems.

Effelsberg (Uwe Bach at Tue Mar 23 11:19:37 2010) Success:

Noto (P. Cassaro at Fri Mar 26 09:43:45 2010) Success:

Jodrell_Bank (alastair gunn at Mon Mar 29 12:42:32 2010) Success:
The Lovell telescope was used. No known problems.